OpenStudy (anonymous):

log 3 36 - 4log 9 30 + 16log 81 (sqrt)15

OpenStudy (anonymous):

\[\log_{3}36-4\log_{9}30+16\log_{81}\sqrt{15} \]

OpenStudy (australopithecus):

you just want to combine these right?

OpenStudy (australopithecus):

or write as one term

OpenStudy (australopithecus):

ugh equation editor isnt working one sec

OpenStudy (australopithecus):

\[\log(a^b) = b \log(a)\]

OpenStudy (australopithecus):

\[-\log(a) = \log(a^{-1}) = \log(\frac{1}{a})\] \[\log_c(a) + \log_c(b) = \log(ab)\] Note however, \[\log_c(a) + \log_d(b) \neq \log(ab)\] as they are different bases

OpenStudy (australopithecus):

can you solve it now?

OpenStudy (australopithecus):

Note that one rule isnt unique to logarithms \[a^{-c} = \frac{1}{a^c}\] \[\frac{1}{a^{-c}} = a^{c}\] One more note, \[a^1 = a\]

OpenStudy (australopithecus):

For the last term in your problem it may be a bit tricky first off note that, \[\sqrt{15} = 15^{\frac{1}{2}}\]

OpenStudy (australopithecus):

so you will need to multiply 16 by 1/2

OpenStudy (australopithecus):

as a general rule I suppose to make things simple, \[\log(a) - \log(b) = \log(\frac{a}{b})\]

OpenStudy (anonymous):

i dont think I have difficulty with the first two terms, but the last one I've been trying to use the change of base formula